OpenStudy (anonymous):

find a area of a rectangle 0.35 yard and 1/5 yards

OpenStudy (anonymous):

if these are the side lengths, simply multiply them to get the area

OpenStudy (anonymous):

so 0.35* 1/5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so can you show me an equation

OpenStudy (anonymous):

A=l*w so: A=0.35*1/5

OpenStudy (anonymous):

will it be 0.07

OpenStudy (anonymous):

Yes, that is correct :)